# Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning

### v3

@@ 3.3 Per-example MIA vulnerability
Although LiRA models $t_{x}$ by a normal distribution, we consider a more general case where the true distribution of $t_{x}$ is of the location-scale family. That is,
$t_{x}=\begin{cases}\mu_{\mathrm{in}}(x)+\sigma_{\mathrm{in}}(x)t&\text{ if }(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\\ \mu_{\mathrm{out}}(x)+\sigma_{\mathrm{out}}(x)t&\text{ if }(x,y_{x})\notin\mathcal{D}_{\mathrm{target}},\end{cases}$ (11)
where $t$ has the standard location and unit scale, and $\mu_{\mathrm{in}}(x),\mu_{\mathrm{out}}(x)$ and $\sigma_{\mathrm{in}}(x),\sigma_{\mathrm{out}}(x)$ are the locations and scales of IN/OUT distributions of $t_{x}$. We assume that the target and shadow datasets have a sufficient number of examples. This allows us to also assume that $\hat{\sigma}(x)=\hat{\sigma}_{\mathrm{in}}(x)=\hat{\sigma}_{\mathrm{out}}(x)$ and $\sigma(x)=\sigma_{\mathrm{in}}(x)=\sigma_{\mathrm{out}}(x)$, where $\hat{\sigma}(x)$ is the standard deviation of $t_{x}$ estimated from shadow models and $\sigma(x)$ is the true scale parameter of $t_{x}$. (See Appendix B for the validity of these assumptions). The following result reduces the LiRA vulnerability to the location and scale parameters of $t_{x}$.
@@
Suppose that the true distribution of $t_{x}$ is of location-scale family with locations $\mu_{\mathrm{in}}(x),\mu_{\mathrm{out}}(x)$ and scale $\sigma(x)$, and that LiRA models $t_{x}$ by $\mathcal{N}(\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}(x))$ and $\mathcal{N}(\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}(x))$. Assume that an attacker has access to the underlying distribution ${\mathbb{D}}$. Then for a large enough number of examples per class and infinitely many shadow models, the LiRA vulnerability of a fixed target example is
$\textsc{tpr}_{\mathrm{LiRA}}(x)=\begin{cases}1-F_{t}\left(F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}(x)-\mu_{\mathrm{out}}(x)}{\sigma(x)}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(x)>\hat{\mu}_{\mathrm{out}}(x)\\ F_{t}\left(F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}(x)-\mu_{\mathrm{out}}(x)}{\sigma(x)}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(x)<\hat{\mu}_{\mathrm{out}}(x),\end{cases}$ (12)
where $F_{t}$ is the cdf of $t$ with the standard location and unit scale.
@@
See Section C.1. ∎
Here we assume that an attacker trains shadow models with the true underlying distribution. However, in real-world settings the precise underlying distribution may not be available for an attacker. We relax this assumption in Appendix B so that the attacker only needs an approximated underlying distribution for the optimal LiRA as in Lemma 1.
Next we focus on the per-example RMIA performance. As in the case of LiRA, we assume that $t_{x}$ and $t_{z}$ follow distributions of the location-scale family. We have
$t_{x}$ =$=\begin{cases}\mu_{x,x}+\sigma_{x,x}t&\text{ if }(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\notin\mathcal{D}_{\mathrm{target}}\\ \mu_{z,x}+\sigma_{z,x}t&\text{ if }(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\in\mathcal{D}_{\mathrm{target}}\end{cases}$ (13)
$t_{z}$ =$=\begin{cases}\mu_{x,z}+\sigma_{x,z}t&\text{ if }(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\notin\mathcal{D}_{\mathrm{target}}\\ \mu_{z,z}+\sigma_{z,z}t&\text{ if }(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\in\mathcal{D}_{\mathrm{target}}.\end{cases}$ (14)
It is important to note that $\mu_{a,b}$ and $\sigma_{a,b}$ denote, respectively, a location and a scale, while previously defined $\hat{\mu}_{a,b}$ and $\hat{\sigma}_{a,b}$ are, respectively, a mean and a standard deviation. As for the analysis of LiRA, we assume that the target and shadow sets have a sufficient number of examples per class, and that $\sigma_{x}=\sigma_{x,x}=\sigma_{z,x}$, $\sigma_{z}=\sigma_{x,z}=\sigma_{z,z}$, $\hat{\sigma}_{x}=\hat{\sigma}_{x,x}=\hat{\sigma}_{z,x}$ and $\hat{\sigma}_{z}=\hat{\sigma}_{x,z}=\hat{\sigma}_{z,z}$, where $\sigma_{x}$ and $\sigma_{z}$ are, respectively, the true scales of $t_{x}$ and $t_{z}$, and $\hat{\sigma}_{x}$ and $\hat{\sigma}_{z}$ are, respectively, standard deviations of $t_{x}$ and $t_{z}$ estimated from shadow models (see Appendix B for the validity of these assumptions).
Lemma 2 (Per-example RMIA vulnerability).
Suppose that the true distributions of $t_{x}$ and $t_{z}$ are of location-scale family with locations $\mu_{x,x},\mu_{z,x},\mu_{x,z},\mu_{z,z}$ and scales $\sigma_{x},\sigma_{z}$, and that RMIA models $t_{x}$ and $t_{z}$ by normal distributions with parameters computed from shadow models. For a large enough number of examples per class and infinitely many shadow models, the RMIA vulnerability of a fixed target example is bounded by
$\textsc{tpr}_{\mathrm{RMIA}}(x)\leq\begin{cases}1-F_{t}\left(F_{t}^{-1}(1-\alpha)-\frac{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[q]}{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[A]}\right)&\text{ if }\ \mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[A]>0\\ F_{t}\left(F_{t}^{-1}(\alpha)-\frac{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[q]}{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[A]}\right)&\text{ if }\ \mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[A]<0,\end{cases}$ (15)
for some constant $\alpha\geq\textsc{fpr}_{\mathrm{RMIA}}(x)$, where
$q$ $=\frac{(\mu_{x,x}-\mu_{z,x})(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})}{\hat{\sigma}_{x}^{2}}-\frac{(\mu_{x,z}-\mu_{z,z})(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}{\hat{\sigma}_{z}^{2}}$ (16)
$A$ $=\frac{\sigma_{x}}{\hat{\sigma}_{x}^{2}}(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\sigma_{z}}{\hat{\sigma}_{z}^{2}}(\hat{\mu}_{x,z}-\hat{\mu}_{z,z}).$ (17)
Proof.
See Section C.2. ∎
Note that here we must assume that the attacker has access to the underlying distribution for the optimal RMIA as the Equations 16 and 17 depend on the parameters computed from shadow models.
3.4 A simplified model of membership inference
@@
We sample vectors on a high dimensional unit sphere and classify them based on inner product with estimated class mean. This model is easier to analyse theoretically than real-world deep learning examples. We generate the data and form the classifiers (which are our target models) as follows:
@@
    For each class, we first sample a true class mean $m_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in\{1,\dots,C\}:m_{i}\perp m_{j}\lor i=j$).
@@
    We sample $2S$ vectors $x_{c}$ for each class. We assume that they are Gaussian distributed around the the true class mean $x_{c}\sim\mathcal{N}(m_{c},s^{2}I)$ where the $s^{2}$ is the in-class variance.
@@
    For each ”target model” we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $r_{c}$.
@@
    The computed mean is used to classify sample $x$ by computing the inner product $\langle x,r_{c}\rangle$ as a metric of similarity.
The attacker has to infer which vectors have been used for training the classifier. Instead of utilising the logits (like in many image classification tasks), the attacker can use the inner products of a point with the cluster means.
@@
Fix a target example $(x,y_{x})$. For the simplified model with arbitrary $C$ and infinitely many shadow models, the per-example LiRA vulnerability is given as
$\textsc{tpr}_{\mathrm{LiRA}}(x)=\Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))+\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}s||x||}\right),$ (18)
where $m_{x}$ is the true mean of class $y_{x}$. In addition, for large $S$ we have
$\log(\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))^{2}+\log\frac{|\langle x,x-m_{x}\rangle|}{||x||s\sqrt{2\pi}}.$ (19)
@@
$\log(\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x))\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))^{2}+\log\frac{||x-m_{x}||}{s\sqrt{2\pi}}.$ (20)
This implies that if $||x-m_{x}||$ is bounded, then the worst-case vulnerability is also bounded. Hence we can have the LiRA vulnerability arbitrarily close to zero in this non-DP setting by simply increasing $S$. Similarly, employing Lemma 2 and the simplified model, we obtain the following upper bound for RMIA performance.
Theorem 4 (Per-example RMIA power-law).
Fix a target example $(x,y_{x})$. For the simplified model with infinitely many shadow models, the per-example RMIA vulnerability is given as
$\textsc{tpr}_{\mathrm{RMIA}}(x)\leq 1-F_{|t|}\left(F_{|t|}^{-1}(1-\alpha)-\frac{\psi(x,C)}{\sqrt{S}s}\right),$ (21)
where $\alpha\geq\textsc{fpr}_{\mathrm{RMIA}}(x)$ and
$\psi(x,C)=\\ \frac{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}\left[2||x-z||^{2}\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}\left[(||x-m_{x}||^{2}+||z-m_{z}||^{2})\mid y_{z}\not=y_{x}\right]}{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}\left[2||x-z||\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}\left[(||x-m_{x}||+||z-m_{z}||)\mid y_{z}\not=y_{x}\right]}.$ (22)
In addition, for large $S$ we have
$\log(\textsc{tpr}_{\mathrm{RMIA}}(x)-\textsc{fpr}_{\mathrm{RMIA}}(x))\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\alpha)^{2}+\log\frac{\psi(x,C)}{\sqrt{\pi/2}}.$ (23)
Proof.
See Section C.4. ∎
As for the LiRA power-law, bounding $||x-z||,||x-m_{x}||$ and $||z-m_{z}||$ will provide a worst-case upper bound for which the power-law holds. Now the following corollaries extend the power-law to the average-case MIA vulnerabilities. We will also empirically validate these results in Section 4.
@@
$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\langle x,x-m_{x}\rangle}{\sqrt{2\pi}||x||s}\right]\right).$ (24)
@@
See Section C.5. ∎
Corollary 6 (Average-case RMIA power-law).
$\log(\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}})\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\alpha)^{2}+\log\left(\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\psi(x,C)}{\sqrt{2\pi}}\right]\right).$ (25)
Proof.
See Section C.6. ∎
@@ 4 Empirical evaluation of MIA vulnerability and dataset properties
In this section, we investigate how different properties of datasets affect the MIA vulnerability. Based on our observations, we propose a method to predict the vulnerability to MIA using these properties.
@@ 4.1 Experimental setup
We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k (Russakovsky et al., 2015). The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section F.1. We optimise the hyperparameters (batch size, learning rate and number of epochs) using the library Optuna (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section E.2). We provide the the code for reproducing the experiments in the supplementary material.
Measuring the uncertainty for tpr The tpr values from the LiRA-based classifier can be seen as maximum likelihood-estimators for the probability of producing true positives among the positive samples. Since we have a finite number of samples for our estimation, it is important to estimate the uncertainty in these estimators. Therefore, when we report the tpr values for a single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals (Clopper & Pearson, 1934). Given tp true positives among p positives, the $1-\alpha$ confidence Clopper-Pearson interval for the tpr is given as
$\begin{split}&B(\alpha/2;\textsc{tp},\textsc{p}-\textsc{tp}+1)<\textsc{tpr}\\ &\textsc{tpr}<B(1-\alpha/2;\textsc{tp}+1,\textsc{p}-\textsc{tp}),\end{split}$ (26)
where $B(q;a,b)$ is the $q$th-quantile of $\text{Beta}(a,b)$ distribution.

### v4

@@ 3.3 Per-example MIA vulnerability
Although LiRA models $t_{\bm{x}}$ by a normal distribution, we consider a more general case where the true distribution of $t_{\bm{x}}$ is of the location-scale family. That is,
$t_{\bm{x}}=\begin{cases}\mu_{\mathrm{in}}(\bm{x})+\sigma_{\mathrm{in}}(\bm{x})Z&\text{ if }(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}\\ \mu_{\mathrm{out}}(\bm{x})+\sigma_{\mathrm{out}}(\bm{x})Z&\text{ if }(\bm{x},y)\notin\mathcal{D}_{\mathrm{target}},\end{cases}$ (7)
where $Z$ has the standard location and unit scale, and $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and $\sigma_{\mathrm{in}}(\bm{x}),\sigma_{\mathrm{out}}(\bm{x})$ are the locations and scales of IN/OUT distributions of $t_{\bm{x}}$. We assume that the target and shadow datasets have a sufficient number of examples. This allows us to also assume that $\hat{\sigma}(\bm{x})=\hat{\sigma}_{\mathrm{in}}(\bm{x})=\hat{\sigma}_{\mathrm{out}}(\bm{x})$ and $\sigma(\bm{x})=\sigma_{\mathrm{in}}(\bm{x})=\sigma_{\mathrm{out}}(\bm{x})$, where $\hat{\sigma}(\bm{x})$ is the standard deviation of $t_{\bm{x}}$ estimated from shadow models and $\sigma(\bm{x})$ is the true scale parameter of $t_{\bm{x}}$. (See Section A.2 for the validity of these assumptions). The following result reduces the LiRA vulnerability to the location and scale parameters of $t_{\bm{x}}$.
@@
Suppose that the true distribution of $t_{\bm{x}}$ is of location-scale family with locations $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and scale $\sigma(\bm{x})$, and that LiRA models $t_{\bm{x}}$ by $\mathcal{N}(\hat{\mu}_{\mathrm{in}}(\bm{x}),\hat{\sigma}(\bm{x})^{2})$ and $\mathcal{N}(\hat{\mu}_{\mathrm{out}}(\bm{x}),\hat{\sigma}(\bm{x})^{2})$. Assume that an attacker has access to the underlying distribution ${\mathbb{D}}$. Then for a large enough number of examples per class and infinitely many shadow models, the LiRA vulnerability of a fixed target example is
$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})=\begin{cases}1-F_{Z}\left(F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(\bm{x})>\hat{\mu}_{\mathrm{out}}(\bm{x})\\ F_{Z}\left(F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(\bm{x})<\hat{\mu}_{\mathrm{out}}(\bm{x}),\end{cases}$ (8)
where $F_{Z}$ is the cdf of $Z$ with the standard location and unit scale, assuming that the inverse of $F_{Z}$ exists.
@@
See Section A.3. ∎
Here we assume that an attacker trains shadow models with the true underlying distribution. However, in real-world settings the precise underlying distribution may not be available for an attacker. We relax this assumption in Section A.4 so that the attacker only needs an approximated underlying distribution for the optimal LiRA as in Lemma 1.
3.4 A simplified model of the optimal membership inference
@@
We sample vectors on a high-dimensional unit sphere and classify them based on inner product with estimated class mean. This model is easier to analyse theoretically than real-world deep learning examples. We generate the data and form the classifiers (which are our target models) as follows:
@@
    For each class, we first sample a true class mean $m_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in\{1,\dots,C\},i\neq j:m_{i}\perp m_{j}$).
@@
    We sample $2S$ vectors $\bm{x}_{c}$ for each class. We assume that they are Gaussian distributed around the the true class mean $\bm{x}_{c}\sim\mathcal{N}(m_{c},\Sigma)$ where the $\Sigma$ is the in-class covariance.
@@
    For each ”target model” we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $\hat{m}_{c}$.
@@
    The computed mean is used to classify sample $\bm{x}$ by computing the inner product $\langle\bm{x},\hat{m}_{c}\rangle$ as a metric of similarity.
The attacker has to infer which vectors have been used for training the classifier. Instead of utilising the logits (like in many image classification tasks), the attacker can use the inner products of a point with the cluster means. Since the inner product score follows a normal distribution, LiRA with infinitely many shadow models is the optimal attack by the Neyman–Pearson lemma (Neyman and Pearson, 1933), which states that the likelihood ratio test is the most powerful test for a given FPR.
@@
Fix a target example $(\bm{x},y)$. For the simplified model with arbitrary $C$ and infinitely many shadow models, the per-example LiRA vulnerability is given as
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S))$ (9)
where $m_{\bm{x}}$ is the true mean of class $y$ and $\xi(S)=O(1/\sqrt{S})$. For large $S$ we have
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (10)
@@
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{||\bm{x}-m_{\bm{x}}||}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}\\ +\log(1+\xi(S)).$ (11)
This implies that if $||\bm{x}-m_{\bm{x}}||$ is bounded, then the worst-case vulnerability is also bounded. Hence we can significantly reduce the MIA vulnerability of all examples in this non-DP setting by simply increasing the number of examples per class.
Now the following corollary extends the power-law to the average-case MIA vulnerability. We will also empirically validate this result in Section 4.
@@
$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}\right]\right).$ (12)
@@
See Section A.6. ∎
@@ 4 Empirical evaluation of MIA vulnerability and dataset properties
In this section, we investigate how different properties of datasets (shots $S$ and number of classes $C$) affect the MIA vulnerability. Based on our observations, we propose a method to predict the vulnerability to MIA using these properties.
@@ 4.1 Experimental setup
Figure 2: Small effect of number of classes $C$ (classes) on MIA vulnerability when attacking a fine-tuned ViT-B Head. The solid line is median and the error bars the min/max bounds for the Clopper-Pearson CIs over 12 seeds ($S=32$).
We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k (Russakovsky et al., 2015). The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section D.1. We optimise the hyperparameters (batch size, learning rate and number of epochs) using the Optuna library (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section C.2). We provide the the code for reproducing the experiments in the supplementary material.